Q.1) Fresh Food has recently carried out a survey of the demand for their seasonal fruits. They observed the following outcomes over the last seven months: Quantity Sold Price Advertising 8500 $ 2 $ 2,800 4700 $ 5 $ 200 5800 $ 3 $ 400 7400 $ 2 $ 500 6200 $ 5 $ 3,200 7300 $ 3 $ 1,800 5600 $ 4 $ 900 Regression (Equation) QSold = 8536.213882 - 835.7223514 Price + 0.592228496 Advertising A. Estimate sales when the price is $50 and advertising is $100,000, stating any assumption that you need to make. B. If the firm charges $50 and cost per unit is $30 but increases advertising to $110,000, what conclusions can you derive in terms of revenues and profits


1
Expert's answer
2021-11-03T11:26:31-0400

Regression (Equation)= Quantities Sold = 8536.213882 - 835.7223514"\times" Price + 0.592228496"\times" Advertising A.


When the price is $50 and advertising is $100,000,sales would be:


= 8536.213882 - 835.7223514"\times" "\$"50 + 0.592228496"\times" "\$" 100,000


=25,972.95"\approx" 25,980


If the firm charges $50 and cost per unit is $30 but increases advertising to $110,000,there would be a subsequent increase in revenues and profit